1. Eliminate a, b and c from, z=ax+by+cxy

2. Eliminate a & b from az+b=a²x+y

3. Solve z(x-y)=x²p-y²q

3. Solve 2p+3q=1



1
Expert's answer
2021-01-21T14:21:07-0500

(i) "z = ax+by+cxy"

partially differentiating with x and y

"\frac{\partial z}{\partial x} = p = a +cy" (1)

"\frac{\partial z}{\partial y} = q = b+cx" (2)


Again differentiating partially first with y, "\frac{\partial^2 z}{\partial x \partial y} = s =c" (3)

From (1), (2) and (3), "z = (p-sy)x+(q-sx)y+sxy = px+qy-sxy"


(ii) "az+b =a^2x+y"

Partially differentiating with x and y

"\frac{\partial z}{\partial x} = p = a" , and "a\frac{\partial z}{\partial y} = 1 \implies pq=1"




(iii) "z(x-y) = x^2p-y^2q"

"\frac{dx}{x^2}=\frac{dy}{-y^2}=\frac{dz}{z(x-y)}"

Taking first two, "\frac{dx}{x^2}=\frac{dy}{-y^2} \implies c = \frac{1}{x}+\frac{1}{y}"


taking last two terms, "\frac{dy}{-y^2}=\frac{dz}{z(x-y)}"

"\frac{dy}{-y^2}=\frac{dz}{z(\frac{y}{cy-1}-y)} \implies \frac{c}{cy-1}dy = \frac{1}{z}dz"


"z = a(cy-1) \implies a = \frac{xz}{y}"

so solution of the equation is, "f((\frac{1}{x}+\frac{1}{y}),\frac{xz}{y}) = 0"





(iv) "2p+3q\n=1"

"\frac{dx}{2}=\frac{dy}{3}=\frac{dz}{1}"

integrating first two, "\frac{1}{2}x + c = \frac{1}{3}y \implies c = \frac{1}{3}y-\frac{1}{2}x"


Integrating last two terms, "\frac{1}{3}y + b = z \implies b = z-\frac{1}{3}y"

So solution is, "f((\frac{1}{3}y-\frac{1}{2}x ),(z-\frac{1}{3}y )) = 0"